OpenStudy (anonymous):

Which expression is equivalent to 3 exponent -2/3 exponent 6? A.)3 exponent 8 B.)3 exponent -8 C.)3 exponent 4 D.)1/3 exponent 4

OpenStudy (anonymous):

D

OpenStudy (anonymous):

If d is 1/(3^4) then its d

OpenStudy (anonymous):

would u like me to rewrite it i the equation editor?

OpenStudy (callisto):

\[(3^{-2/3}) ^{6} = 3^ { (-2/3)(6)} = 3^{-4} = 1/3^4 \]

OpenStudy (anonymous):

no, is it a fraction with 1 all over 3 to the fourth? if yes then its d

OpenStudy (anonymous):

lemme write the problem..hold on

OpenStudy (anonymous):

thats unnecessary

OpenStudy (anonymous):



OpenStudy (anonymous):

\[3^{-2}/3^{6}\]

OpenStudy (anonymous):

how do you get that answer? explain please

OpenStudy (anonymous):

ohhh, now i change my answer to